OpenStudy (anonymous):

x^2+6x? need help don't give straight answers plz

OpenStudy (anonymous):

x^2+6x=0 you mean this?

OpenStudy (anonymous):

yes i'm a little slow at math.

OpenStudy (anonymous):

factor x out first

Parth (parthkohli):

See, take the common factor i.e. x and put that outside the bracket.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x(? + ?) }\)

OpenStudy (anonymous):

x(x+6)=0 and now look which values of x will mach this

Parth (parthkohli):

Yes, now, either x = 0 or x + 6 = 0 x + 6 = 0 x = -6 So, we can infer that x = 0 or -6

OpenStudy (anonymous):

thanks to both of you :)

Parth (parthkohli):

No problem, Amanda. We need keen askers like you on this site :)

OpenStudy (anonymous):

x(x+6)=0 either x = 0 or x+6=0 --> x=-6

OpenStudy (anonymous):

so the answer would be x(x+6) right?

OpenStudy (anonymous):

there are two possible answers x={0,-6}

OpenStudy (anonymous):

thats the only one i have that makes since haha..

OpenStudy (australopithecus):

x^2+6x = 0 first factor it

OpenStudy (inkyvoyd):

If you have around two hours of time, and want to learn completnig the square, go ahead and try to read http://openstudy.com/users/inkyvoyd#/updates/4fab0d4ee4b059b524f6d7b1

OpenStudy (anonymous):

i just need to know if x(x+6) could be an answer?

OpenStudy (australopithecus):

x(x+6) = 0 now lets find the two solutions to this divide both sides by x x(x+6) = 0 Notice that if x = 0 then both sides are zero thus it is one of the solutions if x = - 6 then we have -6(0) = 0 0 = 0 thus -6 is a solution for x

OpenStudy (australopithecus):

sorry if I ruined the question for you but I felt it would be better if I just explained the answers, besides you will see about a million of these problems in the future if you take math in uni or college

OpenStudy (australopithecus):

or if you are already in college or uni other math classes

OpenStudy (anonymous):

i'm in high school and i'm learning college math already so yea lol i'm not good at this.

OpenStudy (australopithecus):

its not hard really just remember to factor and if all else fails use the quadratic formula

OpenStudy (anonymous):

alright thanks alot.

Parth (parthkohli):

Yes, it is x(x + 6)

OpenStudy (australopithecus):

https://www.youtube.com/watch?v=H5AM1bzqCQw

OpenStudy (australopithecus):

this video shows you how to solve the quadratic formula

Parth (parthkohli):

You had to factor right? @amanda132